OpenStudy (anonymous):

I would really appreciate if someone could help me with the following questions below. I really really stink at this.

OpenStudy (anonymous):

Here it is

OpenStudy (anonymous):

\(\Large\rm\color{orange}{Welcome~to~OpenStudy!}\) We can help you but you'll have to figure the answers out on your own because this is a quiz.

OpenStudy (anonymous):

No worries that's what I'm looking for.

OpenStudy (anonymous):

@just_one_last_goodbye @GreatGuy

OpenStudy (anonymous):

@Australopithecus

OpenStudy (phi):

any idea what the little "red marks" mean in your picture?

OpenStudy (anonymous):

What red marks are you referring to?

OpenStudy (phi):

the little red lines on AB and BC

OpenStudy (anonymous):

No clue tbh

OpenStudy (rhystic):

the red ticks indicate that AB and BC are equal in length so if AB = 80 you have your answer. This is standard for any problem so commit it to memory; if there are more then one set of lines with same lengths they are indicated with two tick marks each for second set etc.

OpenStudy (phi):

oh. When you see a "hash mark" like those, it means the two line segments are the same length

OpenStudy (phi):

so you should think: from A to B is some distance. and from A to C is twice the distance from A to B

OpenStudy (anonymous):

Okay so it would be 160

OpenStudy (phi):

now you need to read what they told you. what did they tell you ?

OpenStudy (phi):

do you see the question: if AC is 80 ... that means the distance from A to C is 80 (inches, miles, whatever) they want to know A to B we know AB + BC = AC (the two halves add up to the whole)

OpenStudy (anonymous):

ahh 40

OpenStudy (phi):

yes.

OpenStudy (phi):

For Q6, if AB is 28, then BC is what ?

OpenStudy (phi):

it is the same picture and the same idea

OpenStudy (anonymous):

56

OpenStudy (phi):

AB and BC are the same size (that is what the red lines mean)

OpenStudy (anonymous):

So they are both 28?

OpenStudy (phi):

yes AB and BC are both 28

OpenStudy (phi):

they say BC is also 3x+7 so if BC =28 and BC= 3x +7 we can say 3x+7= 28 can you "solve for x" ?